### v2

# DirectMultiStep: Direct Route Generation for Multistep Retrosynthesis

Abstract
Traditional computer-aided synthesis planning (CASP) methods rely on iterative single-step predictions, leading to exponential search space growth that limits efficiency and scalability. We introduce a series of transformer-based models, utilizing mixture of experts approach, that directly generate multistep synthetic routes as a single string by conditionally predicting each molecule based on all preceding ones. Our models can accommodate specific conditions such as the desired number of steps and starting materials, with the top-performing DMS-Flex (Duo) surpassing state-of-the-art methods on the PaRoutes dataset with a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. It also successfully predicts routes for FDA-approved drugs not included in the training data, showcasing its generalization capabilities. While the current suboptimal diversity of the training set may impact performance on less common reaction types, our multistep-first approach presents a promising direction towards fully automated retrosynthetic planning.
1 Introduction
Finding the most efficient route to a desired chemical compound is a daily challenge for synthetic organic chemists in both fundamental research and drug discovery. Route efficiency is determined by various factors, some of which can be objectively assessed, such as overall yield (not all chemical reactions have 100% conversion rate) and enantiomeric excess (in case of chiral compounds), where higher values are always preferred. Other factors, such as atom efficiency (minimization of byproducts) and availability (cost) of starting materials, are more case-dependent. A used chemical reactant is considered waste unless it can be repurposed as a reactant in a different process. Similarly, the choice of starting materials depends on factors such as budget, logistics, and the availability of specific equipment. It’s worth noting that many commercially available compounds can be synthesized from other commercially available compounds, adding another layer of complexity to the decision-making process.
Algorithmic frameworks for identifying synthetic routes were envisioned by Vleduts [1] and further formalized by Elias James Corey (subject of the 1990 Nobel Prize in Chemistry) into what is now known as retrosynthetic analysis. This framework begins with identifying atoms that would serve as reaction centers. Disconnecting bonds between these centers results in the formation of hypothetical fragments (called synthons) from which a precursor molecule can be created. This mapping from synthons to actual molecules is one-to-many because there is usually more than one functional group that could participate in any given type of reaction. Importantly, meticulous application of Corey’s framework (i.e., systematically breaking small subsets of bonds) will eventually lead to commercially available starting materials. The algorithmic nature of this process allowed Corey to envision automating these rules to create Computer-Aided Synthesis Planning (CASP) as early as 1969 [2].
Recent advancements in data science and machine learning (ML) methods have led to a surge of interest in developing CASP methods [3, 4]. The vast majority of existing methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] are designed to automate single-step retrosynthetic (SSR) analysis, i.e. predicting a list of compounds from which a target product could be made in one step. A full multistep route can be determined by iteratively applying SSR methods to the identified precursors until a termination condition is satisfied (e.g., identifying reactants in the database of commercially available compounds). Notably, because each SSR method call creates a list of candidates, iterative application of these methods generates an exponentially growing search space. Graph traversal algorithms such as Monte Carlo Tree Search (MCTS), Depth-First Proof Number (DFPN) search, and A-star (Retro*) have been applied to efficiently traverse this exponential search space [37, 38, 39]. More sophisticated search designs include utilization of hyper-graph [40], graph-based neural networks [41], reinforcement learning [42], or incorporation of previous reactions as context for single-step predictions [43].
Evaluation of performance of these SSR methods on full route prediction has been limited by the scarcity of open-source datasets containing valid multistep routes. The creation of PaRoutes [44], an open-source dataset containing 450k multistep routes (163k excluding duplicates and permutations, along with two test sets of 10k routes each) extracted from the United States Patent and Trademark Office (USPTO), marked a major development. Notably, even state-of-the-art SSR models [7, 45] combined with advanced search algorithms (MCTS, DFPN, Retro*) correctly identify multistep routes (Top-1 accuracy) for only 17% and 10% of the target compounds in n1 and n5 test sets, respectively. Such performance could be rationalized by recognizing that retrosynthesis is inherently a multistep problem: the optimal choice of reaction to make a compound depends on subsequent steps in the synthesis. For example, a common pattern in multistep routes includes: (1) protection of certain functional groups, (2) the desired transformation of unprotected groups, and (3) deprotection of the protected groups. An SSR method applied to the reactant of reaction (2) may output numerous candidate precursors; however, knowing that protective groups are removed in subsequent steps dramatically changes the probability distribution over those precursors.
In this work, we propose a novel approach for direct prediction of multistep routes, bypassing the need for single-step models and sophisticated exponential graph traversal algorithms. We begin with a brief overview of the methodology and evaluation metrics. We proceed by demonstrating and discussing performance of DMS models on diverse n1 and n5 evaluation sets (10 000 routes each). We demonstrate further generalizability by successfully predicting experimental synthetic routes for FDA-approved drugs. We conclude by discussing the limitations of our approach. A more elaborate description of model architectures and rationale for certain design choices is given at the end of the paper.
2 Methodology Overview
2.1 Definitions
Target compound (blue in Fig. 1) is the final (desired) product of the multistep synthesis tree. Starting material, SM (red in Fig. 1) is a compound for which no further precursors need to be identified, i.e., a leaf of the synthesis tree. The number of steps is the largest number of reactions from SM to the target compound (tree height).
2.2 Route Representation
Routes are represented as recursive dictionaries (Fig. 1) containing the SMILES representation of the molecule and a list of other dictionaries, containing either starting materials or trees leading to precursors of the current node. Removing space and newline characters creates a string representation. Our models predict that string by taking SMILES of the target compound, (optionally) starting material, and (optionally) number of steps as input.
[…]
Figure 1: The workflow of DirectMultiStep. (a) The SMILES representation of the target compound (blue), (optional) starting material (red), and the number of steps are tokenized, concatenated, and fed into our transformer model. The model predicts a string representation of the multistep synthesis tree. Spaces are added for clarity, and indentations indicate the levels in the synthesis route (tree). (b) Molecular structures corresponding to the target compound (blue), (optional) starting material (red), and the predicted synthesis tree with structures of all molecules.
2.3 Evaluation Metrics and Post-Processing
Evaluation is performed on n1 and n5 (10 000 routes each) evaluation sets from PaRoutes [44], which have been created by selecting the most diverse routes from the USPTO. Fig. 2 shows the distribution of route lengths in training set, n1, and n5. The distribution in both n1 and n5 is shifted towards longer routes; for example, the fraction of routes with 4 steps or more is 15% for training set, 27% for n1 and 51% for n5.
Figure 2: Distribution of the relative frequencies of route lengths (in terms of number of steps) in the training dataset before augmentation with permutations (163 689 routes, black), n1 test set (10 000 routes, blue), and n5 test set (10 000 routes, purple). Distribution is split into routes shorter (left subplot) and longer than 6 steps (right subplot).
The structure of target compound, starting material (optionally, the deepest leaf of the correct route is chosen during evaluation), and number of steps is passed to the transformer encoder. A beam search with a width of 50 is employed to generate 50 candidate routes by sampling next tokens from the decoder. Within beam search, during the prediction of each token the cumulative log likelihoods are normalized by the sequence length at the time the new token is sampled. This adjustment ensures that shorter routes with simpler compounds are not disproportionately favored while maintaining a balance between route length and overall sequence probability. After beam search finds end-of-sequence tokens for all 50 candidates, predicted routes are checked for validity of all SMILES strings (which are canonicalized), the presence of starting materials in the stock set (following procedure of PaRoutes taken to be the set of all leaves in n1 or n5, and absence of repetitions (which can be present as permutations of other routes). Remaining routes are compared against the true (experimental) route. The reproduction of experimental routes is presented as Top-$K$ accuracy, i.e. presence of correct route ranked $K$ or lower. Importantly, the Top-$K$ accuracy is evaluated against all 10 000 targets, not just the ones for which routes are found.
2.4 Models Presented
As a baseline, we train a conventional transformer-based sequence-to-sequence model with 6 encoder and 6 decoder layers (DMS-Flash) with structures of starting materials provided during training. To achieve comparable performance without starting material as a hint, a significantly larger model with 12 encoder and 36 decoder layers is needed (DMS-Deep). We also present the first, to our knowledge, application of mixture-of-experts (MoE [46, 47, 48]) architecture to a chemical problem. To create MoE model, one replaces the 2-layer MLP that follows every attention block with an ensemble of 2-layer MLPs (called experts). A preceding router determines (for each input) the $n$-most relevant (active) experts, and passes the input through each of them. Output of each active expert is added with weights determined by the router. Notably, the number of experts active during inference does not have to be equal to the number of experts active during training, although the latter tends to be optimal. We present two MoE models, the DMS-Flex (trained with SM) and DMS-Wide (trained without SM). The relevant architectural details of models are given in Table 1.
Table 1: Specifications of architecture for different model variants.
DMS Model Encoder Decoder Hidden FF Total Active Experts
Variant Layers Layers Dimension Multipliera Experts Training Generation
Flash 6 6 256 3 1 1 1
Explorer 6 6 256 3 3 2 2
Explorer XL 6 24 256 3 3 2 2
Flex (Mono) 6 6 256 3 3 2 1
Flex (Duo) 6 6 256 3 3 2 2
Deep 12 36 256 3 1 1 1
Wide 12 12 256 3 3 2 2
a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
3 Results and Discussion
3.1 Top-K Accuracy on Evaluation Sets
3.1.1 Vanilla Transformers
Top-$K$ accuracy is shown in Table 2 for set-n1 and Table 3 for set-n5. The smallest DMS-Flash model (which receives structure of SM during generation) achieves state-of-the-art performance, surpassing all previous methods in all Top-$K$ accuracies, with the most significant increase in Top-1 accuracy: 2.4x on n1 and 3.7x on n5. To achieve comparable performance without providing the structure of the SM, a significantly larger model DMS-Deep is needed. This establishes a trade-off between size of the model (and thus generation speed) and the amount of information it receives. The limitations of the necessity to provide the structure of the SM are discussed in Section 3.1.3. Quite remarkably, a Flash model trained with structures of SM may generate routes even when SMs are not provided (Tables 2-3). While that results in a 20% drop in Top-1 accuracy, it is significantly better than a model with the same architecture that is trained without SM (Tables S6-S7). In other words, providing structure of a correct starting material helps the model learn better and improves performance even if the starting material is not used during evaluation.
Table 2: Top-$K$ accuracy on route test set-n1 (10 000 routes).
Method Top-1 Top-5 Top-10
MCTSa 0.17 0.46 0.49
Retro*a 0.15 0.41 0.45
DFPNa 0.11 0.17 0.17
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.41 0.50 0.54 0.56 0.58 0.60
Flash 10M 0.32 0.39 0.42 0.44 0.45 0.47
Explorer 19M 0.36 0.45 0.48 0.51 0.52 0.55
Explorer 19M 0.27 0.33 0.36 0.37 0.39 0.40
Explorer-XL 50M 0.32 0.40 0.44 0.45 0.46 0.48
Flex (Mono) 19M 0.32 0.40 0.43 0.44 0.45 0.48
Flex (Duo) 19M 0.43 0.52 0.55 0.57 0.58 0.61
Flex (Duo) 19M 0.34 0.41 0.44 0.45 0.46 0.47
Deep 41M 0.36 0.44 0.47 0.49 0.50 0.52
Wide 38M 0.38 0.46 0.49 0.50 0.51 0.53
aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
Table 3: Top-$K$ accuracy on route test set-n5 (10 000 routes).
Method Top-1 Top-5 Top-10
MCTSa 0.10 0.28 0.33
Retro*a 0.10 0.30 0.36
DFPNa 0.05 0.07 0.07
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.37 0.44 0.47 0.49 0.51 0.53
Flash 10M 0.30 0.36 0.39 0.40 0.41 0.43
Explorer 19M 0.33 0.40 0.43 0.45 0.46 0.49
Explorer 19M 0.25 0.31 0.34 0.35 0.36 0.37
Explorer-XL 50M 0.31 0.38 0.41 0.42 0.43 0.45
Flex (Mono) 19M 0.28 0.34 0.36 0.38 0.39 0.41
Flex (Duo) 19M 0.39 0.46 0.49 0.51 0.52 0.54
Flex (Duo) 19M 0.32 0.38 0.40 0.42 0.42 0.44
Deep 41M 0.33 0.40 0.43 0.45 0.46 0.48
Wide 38M 0.35 0.42 0.44 0.46 0.47 0.49
aThese data are collected from the 2.0 version of PaRoutes in their GitHub repository [44].
Remarkably, the Top-2 accuracy of DMS-Flash and the Top-4 accuracy of DMS-Deep exceed the Top-10 accuracy of all search methods that utilize iterative application of single-step models, highlighting the advantages of the multistep-first approach. To understand this performance difference, it helps to notice that iterative application of single-step calls results in a breadth-first search-like exploration of a node, as all reactants leading to a compound are generated at the same time. Our approach, however, is a depth-first search-like: the whole left sub-tree of target compound in Fig. 1 is predicted before the immediate right child. This provides an extra context for the prediction of the second precursor to the target compound (or any other right child), which reduces the search space and finds more relevant compounds faster.
A few words on the relevance of Top-$K$ accuracy as the primary evaluation metric are in order. One might question whether the number of targets for which valid (i.e., having valid SMILES) routes are found (or solvability rate if all SMs are in a given stock set) and the average length of routes predicted should be considered. It is essential to realize that even if you ensure that every single-step transformation corresponds to a known reaction by template matching, the route may still be experimentally not feasible simply because atoms that are unchanged (i.e., the ones ignored by the template) contribute to the energies of the reacting orbitals through inductive and resonance effects, thereby changing the energetic favorability of the reaction. As a result, short of performing highly-accurate ab initio calculations of energies (after corresponding geometry optimization) with all relevant solvation effects and considering thermal corrections, there is no computational metric that would assess experimental feasibility of the predicted route. In addition, one could argue that the ability to discover an experimentally verified route is a prerequisite for finding shorter, but still feasible, routes. Therefore, Top-$K$ accuracy is the closest proxy to route quality. Additionally, the ability to predict correct routes with better ranks is a significant improvement because it reduces the number of candidates to be passed through the aforementioned highly expensive energy calculations. Finally, one could argue that single-step methods do not provide inherent ranking of candidates, so the performance of search methods may depend on the choice of ranking metric. However, that is yet another benefit of our approach: there is no need in devising another algorithm or model for route ranking, the beam probabilities are sufficiently accurate.
The superior performance of DMS models does not incur additional computational cost. A search with a beam size of 50 takes 3-16 seconds (depending on the number of steps) with a DMS-Flash or a DMS-Flex model on a single GPU (NVIDIA M4000), which is comparable to the time required to find the first successful route using previous methods (7-50 seconds [44]).
Figure 3: Distribution of Top-1 and Top-10 accuracy of predictions with DMS-Flex (Duo) on test sets n1 and n5. There is only one route with length 10 in n1, and DMS-Flex (Duo) does not predict it correctly. That route is reproduced by splitting it in half, as shown in Fig. 4
3.1.2 MoE Transformers
The DMS-Flex (Flexible), a mixture-of-experts model trained with two out of three experts active, provides the largest performance improvement when evaluated with two experts active (Duo mode): 2.5x increase in Top-1 on n1 and 3.9x on n5. Evaluating DMS-Flex with one expert active (Mono) mode worsens the performance by 30%; given that the number of valid routes (Tables S1-S2) does not change significantly, one could consider the Mono mode as a source of diversity in predictions. Notably, even though the DMS-Flex is trained with SM, it can be evaluated without providing SM (empty padding tokens are fed into the encoder instead) which still results in 34% Top-1 accuracy and 47+Top-10. Evaluating DMS-Flex without providing the number of steps as an input drops performance significantly, but a different model with the same architecture trained without number of steps as an input, DMS-Explorer (Fig. S1 shows the distribution of route lengths predicted by Explorer relative to the length of an experimental route), results in only a 8% drop in accuracy.
Using a larger DMS-Wide model trained without SM provides comparable performance to that of DMS-Deep. An important advantage of DMS-Wide is that tokens passed through different active experts can be processed on a GPU in parallel, which enables faster predictions. In addition, DMS-Wide has fewer layers, so fewer consecutive calculations have to be performed; in practice, doubling the number of encoder layers results in doubling of generation time. One might also conclude that similar performance of DMS-Deep and DMS-Wide, which have comparable parameter counts, implies that the expressive power mostly depends on the number of parameters, and not on the extra non-linearities provided by larger number of layers.
Our experimentation with the increase in the number of experts, the number of encoder/decoder layers, or the feed-forward multiplier evaluated on small subsets of n1 or n5, show no significant improvement in performance (Tables S4-S7), which most likely means that the 20 M (with SM) or 40 M (without SM) are optimal given the current size of the training set.
Fig. 3 shows the distribution of Top-1 and Top-10 accuracy over different route lengths on both n1 and n5. Given that 85% of routes in the training partition have 4 or fewer steps Fig. 2), one would expect the accuracy to decrease dramatically with increasing route length. However, the performance on routes with 5-8 steps is comparable to that of shorter routes, and the performance on 9-step routes is even comparable to that of 2-step routes. The exceptional accuracy on 9-step routes could be rationalized by the fact that their relative abundance in n1 and n5 does not exceed the relative abundance in the training partition (unlike for routes with length 3-7, as seen in Fig. 3). Top-1 and Top-10 accuracies of other DMS models for different route lengths are given in Tables S8-S11.
[…]
Figure 4: Separation of a 10-step route from set-n1 into two 5-step routes. (a) Correct prediction from DMS-Flex (Duo) for the first half of the 10-step route with starting material information (red). (b) Correct prediction from DMS-Flex (Duo) for the second half of the 10-step route with starting material information.
3.1.3 Use Cases for Different DMS Model Variants
The requirement to provide the starting material structure as input to the model might, in theory, seem as a limitation, limiting the relevance of DMS-Flash/Flex models. However, in practice, a chemist often can know the structure of a desired starting material because Corey’s retrosynthetic framework allows finding at least one route to starting materials by performing one functional group transformation at a time. Such deterministically found routes will often be too long to be efficient in terms of yield (overall yield is the product of yields of individual steps). The art of organic synthesis lies in finding ways to perform several transformations simultaneously using a minimal number of protecting groups. As such, we believe our SM-based models provide value as an assistant to this task: finding better ways to get to the target compound from a starting material. The requirement to specify the number of steps as an input becomes a useful option in this context. Moreover, if the target compound is chiral (i.e., it has a mirror image which cannot be superimposed), even during manual retrosynthetic planning, chemists check if chiral centers can be obtained from chiral starting materials (such as naturally occurring amino acids) because developing reactions that create chiral centers in the desired conformation remains an active research area (and subject of 2001 and 2021 Nobel Prizes in Chemistry). An example of such a route is shown in Fig. 7.
In addition, one can use more computationally expensive DMS-Wide/Deep models to discover potential starting materials, and then use those structures as inputs to DMS-Flash/Flex to generate suggestions for each starting material of interest.
To lift the necessity to specify desired route length, a user also has a choice between running Flex/Wide models several times with different desired route lengths or simply running Explorer-XL once. Of course, ideally one would prefer a single model that predicts the whole route with just from the structure of the target compound, without number of steps or starting material provided as an extra input. We believe better performance could be attainable in the future once bigger datasets become available. Finding ways to improve performance even with current dataset constitutes an important area of further research.
3.2 Separation of Long Routes into Shorter Routes
As seen in Fig. 3, DMS-Flex (Duo) does not find the single 10-step route in the n1 set. However, this route can be correctly reproduced with rank 1 if split into two 5-step routes and the intermediate compound is provided as the starting material for the first 5 reactions (Fig. 4) and a target compound for the last 5 reactions (Fig. 4). While ideally the necessity to provide an intermediate compound should be avoided, this resembles the bidirectional search technique commonly employed in manual retrosynthetic analysis [49]. In contrast, models trained without SM, such as DMS-Deep and DMS-Wide, can replicate the second 5-step sequence in Fig. 4 but not the first (Fig. 4).
[…]
Figure 5: Literature routes for Vonoprazan and Mitapivat correctly reproduced by DMS-Flex (Duo). Target compounds are in blue, starting materials that are given as inputs are colored in red. Ranks denote the rank of this route when the specified starting material is provided. (a) First literature route for Vonoprazan. The model predicts correctly no matter which starting material is given. (b) Second literature route for Vonoprazan. The model predicts the route correctly only when an immediate precursor to Vonoprazan is given as the target compound. (c) First literature route for Mitapivat. (d) Second literature route for Mitapivat.
3.3 Retrosynthetic planning of pharmaceutical compounds
For another generalizability assessment of our model, we test it on three FDA-approved drugs: Vonoprazan, Mitapivat, and Daridorexant, which were used for evaluation by Xiong, et al. [50]. These drugs and their intermediates are absent from the training set. Vonoprazan, a potassium competitive acid blocker for Helicobacter pylori infections, was initially proposed with a 2-step route (Fig. 5[51]. However, one of its starting materials is unstable, and the final reaction leads to significant byproducts. Subsequently, an improved 4-step route (Fig. 5) was introduced [52]. Mitapivat, a pyruvate kinase activator for treating hemolytic anemia, has been associated with both a 5-step route (Fig. 5) and a 3-step route (Fig. 5[53, 54]. Daridorexant, an orexin receptor antagonist for adult insomnia, is linked to a 4-step route (Fig. 7[55].
Both vanilla transformers (DMS-Flash) and MoE models (DMS-Flex Duo) correctly reproduce Vonoprazan’s first literature route (Fig. 5) regardless of the starting material provided, with high ranks. However, DMS-Flash struggles to find the first transformation (counting from the target compound) in the second route. This inability could be attributed to the fact that the reduction of amide carbonyl (the step leading to Vonoprazan in second route) is less frequently represented in the PaRoutes dataset than the much more common reductive amination (the step leading to Vonoprazan in first route). As a result, most routes predicted by DMS-Flash start with a reductive amination as in Fig. 5. However, if we provide the precursor to Vonoprazan in the second route as the target compound, the model correctly reproduces the remaining route (Fig. 5). This underscores that the performance of our models relies upon sufficient representation of transformations in the training set. This limitation, however, should naturally disappear once larger multistep datasets become available.
Inability of DMS-Flash to suggest reduction of amide carbonyl in Fig. 5 was the motivation for the creation of mixture-of-experts (MoE) models. We conjectured that different experts might specialize in different classes of transformations, and even if one expert dominates the generations, one could still manually force the output from a different expert. Fortunately, no such manual interventions were needed as DMS-Flex (Duo) predicted the route almost identical to the second transformation as the Top-1 candidate (Fig. 6). The only difference from the correct route is the order of transformations: in the correct route the tosyl group is introduced after amidation with methylamine, not before.
Figure 6: A variation of second Vonoprazan route predicted by DMS-Flex (Duo). The correct route is shown in Fig. 5 and differs by the order of the last steps: tosyl group is introduced after amidation, not before.
Figure 7: Daridorexant literature multistep routes and predictions from DMS-Flex (Duo). The compounds that are missing from the prediction are in dotted boxes.
DMS-Flex (Duo) correctly reproduces both routes for Mitapivat (Fig. 5 and Fig. 5). In contrast, the route for Daridorexant (Fig. 7) is only partially predicted: our model does not split the right child of the root node into the two compounds in the dotted boxes. All other intermediates and starting materials are predicted correctly. The reason for this behavior is not well understood, but it can be rationalized by looking at the distribution of the number of leaves at root nodes: 73% of root nodes (target compounds) have at least 1 leaf (see Fig. S3). This dependence on the distributions of the training set comprises an important limitation of our models; however, it should also be resolved once bigger datasets become available. Admittedly, this issue would not occur with current methods based on single-step prediction because the tree search framework would not halt at compounds outside the stock compound set before reaching the maximum iterations. Ensuring that multistep-first models reproduce that aspect of SSR methods constitutes a direction for future research.
This example, however, does not imply that the model struggles to predict routes with convergent branches. If one defines a convergent route as the one that has at least one node (compound) with at least two children that are not leaves, and repeats the analysis, no significant differences in Top-K accuracies (Fig. S5-S6) or in the number of routes found (Fig. S7-S8) is observed despite significant variation in the number of routes that are convergent among training routes, n1, and n5 (Fig. S3).
4 Conclusion
In this work, we introduce the first-of-its-kind inherently multistep approach for multistep retrosynthesis prediction. By predicting the entire route as a single string, our approach bypasses the need for iterative single-step predictions and exponential search space traversal. We present a family of transformer-based DirectMultiStep (DMS) models which can predict correct routes just from the structure of the target compound, but can also accommodate specific conditions such as the desired number of steps and starting materials, enabling efficient route planning.
Our flagman DMS-Flex (Duo) model, incorporating mixture-of-experts approach, outperforms state-of-the-art methods on the PaRoutes dataset, achieving a 2.5x improvement in Top-1 accuracy on the n1 test set and a 3.9x improvement on the n5 test set. We demonstrate the quality of predicted routes by testing our models on several FDA-approved drugs. Those predictions also highlight limitations caused by the suboptimal diversity of the current training set, particularly for less common reaction types. Those limitations are partially mitigated by employing mixture-of-experts techniques, making them a promising candidate for further research in retrosynthesis prediction and applications of generative models to chemistry in general.
We believe our work demonstrates the superiority of multistep-first approach to full route prediction and inspires further research into methods that reduce the exponential complexity of iterative application of single-step based methods. We are grateful to the authors of PaRoutes for the development and release of the first multistep prediction benchmark, and we hope more benchmarks based on experimentally verified routes become available. Future research may focus on improvement of performance on longer routes (e.g. eliminating the necessity to split longer routes) and (or) incorporation of additional constraints (such as cost) on the route prediction. Overall, this work demonstrates the potential of transformer-based models for efficient and scalable computer-aided synthesis planning. We release the code used for training and route prediction as a public GitHub Repo under MIT License. Our models are also available with a Web GUI at models.batistalab.com.
5 More Computational Details
5.1 Data Curation and Preparation
A second version of the PaRoutes [44] dataset containing 450k routes is stripped of all metadata and stored as recursive dictionaries representing multistep routes. The two evaluation sets, n1 and n5, are processed similarly. A training partition is created by removing all repetitions (some routes were represented multiple times with different metadata) and permutations (swapping left and right subtrees) of routes in n1 and n5 from the full dataset, resulting in 163k routes. The training dataset is augmented by adding 2 permutations for each route. To train the DMS model which takes starting material as an input, we find all starting materials (leaves) for each tree and store a combination of the target compound with each starting material as a separate entry. As a result, DMS with SM is trained on 1 340 243 inputs, and the DMS without SM on 432 684 inputs. All SMILES strings are tokenized by treating each character as a token, and the string representation of the multistep route is tokenized similarly, treating delimiters of the tree (‘smiles’, ‘children’, ,, [, ], {, and }) as separate tokens. The final vocabulary size is 52 (including start, end, and padding tokens), the largest multistep route has 1074 tokens, while the largest target compound and starting material have 145 and 135 tokens, respectively.
5.2 Model Architecture
We present a family of DirectMultiStep (DMS) models with varying parameter counts and architectural details. DMS-Flash and DMS-Deep the classical transformer encoder-decoder architecture [56], whereas DMS-Flex and DMS-Wide incorporate sparsely-gated mixture-of-experts layers [46, 47, 48]. Major parameters choices are given in model card (Table  1). All models use GeLU activations [57].
Intermediate experiments with evaluations on subsets of n1 and n5 with 50 and 500 routes reveal no benefit in increasing the number of hidden dimensions to 512 (Tables S4-S5), which could be rationalized by noticing that 256 is already almost 5x the vocabulary size. Although a feed-forward multiplier of four is more common in NLP models, we find no significant improvement in performance as compared to the feed-forward multiplier of three (Tables S6-S7). In addition, the number of encoder layers does not have to be the same as the number of decoder layers, and models with 6 encoder and 12 decoder layers (trained with SM) performed almost as good as the models with 12 encoder and decoder layers. For MoE models, we find no significant improvement upon increase of number of experts to four, and models trained with two active experts outperform those trained with one.
5.3 Training
During training, a 10% dropout rate is used in embedding, attention, and feed-forward layers. The AdamW optimizer [58, 2019adamw] is used with a learning rate scheduler that warms up from 0 to $3.5\times 10^{-4}$ for DMS-Flex and DMS-Wide, $3\times 10^{-4}$ for DMS-Flash, and $1\times 10^{-4}$ for DMS-Deep during the first 10% of training steps and undergoes cosine decay to $10\%$ of the maximal learning rate over 20 epochs. The optimal initial learning rate for each model was chosen by starting from 5e-4 and monitoring for gradient explosion (in which case the learning rate was decreased by 0.5e-4). Gradients are clipped to 1.0. In addition, each token in the input to the encoders has a 5% chance of being masked. During hyperparameter optimization, the model was validated on a 5% split from the training routes (which exclude n1 and n5 routes).
6 Data and Software Availability
The reaction route dataset is from the 2.0 version of PaRoutes in their GitHub repository [44], available under the Apache License 2.0. Code to process dataset, implementation of model architecture, code for training, generation, and evaluation are available under MIT License at https://github.com/batistagroup/DirectMultiStep. A web interface is also available at models.batistalab.com.
Author Contributions
YS, HL, and AM contributed equally to this work. YS conceived the idea. YS and HL conducted preliminary model training and testing. AM introduced the mixture of experts (MoE) approach and carried out additional model training and testing. VSB provided guidance for the research. All authors participated in writing and reviewing the manuscript.
Conflict of Interest
The authors declare no conflict of interest.
7 Acknowledgement
The authors acknowledge a generous allocation of high-performance computing time from NERSC. The development of the methodology was supported by the NSF CCI grant (VSB, Award Number 2124511).
Supporting Information
Figure S1: Lengths of valid routes predicted by DMS-Explorer for targets in n1 and n5 compared to the length of the correct (experimental) route. Left: distribution of differences in lengths for all predictions for all targets. Right: distribution of average (among all predictions for a single target) differences.
Figure S2: Lengths of valid routes predicted by DMS-Flex (Duo) for targets in n1 and n5 compared to the length of the correct (experimental) route. Left: distribution of differences in lengths for all predictions for all targets. Right: distribution of average (among all predictions for a single target) differences.
Figure S3: Distribution of the relative frequencies for the number of leaves (nodes with no children) at root node.
Table S1: Number of predicted routes in n1 that pass post-processing filters.
Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
Flash 9998 9942 8814 8725
Flashe 9994 7660
Explorer 9997 9923 8681 8555
Explorere 9995 7418
Explorer-XLe 9998 8008
Flex (Mono) 9941 9545 8338 7948
Flex (Duo) 9995 9866 8828 8649
Flex (Duo)e 9994 7692
Deepe 9998 7957
Widee 9996 8089
aAt least one route with all valid SMILES (including target, intermediates, and SMs).
bAt least one route with the correct conditioned SMs (and with all valid SMILES).
cAt least one route with all SMs in the n1 stock set (and with all valid SMILES).
dAt least one route with the correct conditioned SMs and with all SMs in the n1 stock set (and with all valid SMILES).
eSMs not provided during inference.
fBlank indicates not relevant (models without SM conditions).
Table S2: Number of predicted routes in n5 that pass post-processing filters.
Model Valida Correct SMb,f SMs In Stockc Correct and In Stockd,f
Flash 9991 9888 8646 8487
Flashe 9991 7576
Explorer 9997 9903 8439 8281
Explorere 9998 7227
Explorer-XLe 10000 7904
Flex (Mono) 9932 9374 8037 7528
Flex (Duo) 9993 9808 8627 8382
Flex (Duo)e 9990 7605
Deepe 9997 7830
Widee 9997 7950
aAt least one route with all valid SMILES (including target, intermediates, and SMs).
bAt least one route with the correct conditioned SMs (and with all valid SMILES).
cAt least one route with all SMs in the n5 stock set (and with all valid SMILES).
dAt least one route with the correct conditioned SMs and with all SMs in the n5 stock set (and with all valid SMILES).
eSMs not provided during inference.
fBlank indicates not relevant (models without SM conditions).
Table S3: Specifications of architecture for different model variants.
DMS Model Encoder Decoder Hidden FF Total Active Experts
Variant Layers Layers Dimension Multipliera Experts Training Generation
Van-19 6 6 256 3 1 1 1
Van-60 8 8 512 4 1 1 1
MoE-1/3 6 6 256 3 3 1 1
a an integer by which the hidden dimension is multiplied in feed-forward layers after attention.
Figure S4: Fraction of routes (grouped by route length) that are convergent in training set, n1, and n5. A route is considered convergent if any node has at least two children that are not leaves. Overall, convergent routes constitute 9.4% of the training routes, 5.6% of n1, and 10.9% of n5.
Figure S5: Top-K accuracy on the n1 test set for different models for (a) all routes in n1, (b) convergent routes in n1, and (c) non-convergent routes in n1. A route is considered convergent if any node has at least two children that are not leaves.
Figure S6: Top-K accuracy on the n5 test set for different models for (a) all routes in n5, (b) convergent routes in n5, and (c) non-convergent routes in n5. A route is considered convergent if any node has at least two children that are not leaves.
Figure S7: Distribution of the average number of routes for targets in n1 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Figure S8: Distribution of the average number of routes for targets in n5 test set (a-c) after checking for the validity of smiles, (d-f) after canonicalization of SMILES and removing repetitions (considering permutations of children), and (g-i) after filtering for routes that have all reactants in stock. Distribution is shown separately for all targets (a, d, g), for targets for which an experimental route is convergent (b, e, h) and for targets for which an experimenal route is non-convergent (c, f, i). A route is considered convergent if any node has at least two children that are not leaves.
Table S4: Top-$K$ accuracy on route test set-n1 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.44 0.53 0.57 0.59 0.60 0.63
MoE-1/3 38M 0.40 0.47 0.50 0.51 0.52 0.53
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S5: Top-$K$ accuracy on route test set-n5 (10 000 routes).
DMS SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Van-19 19M 0.39 0.47 0.50 0.52 0.53 0.56
MoE-1/3 38M 0.36 0.42 0.45 0.46 0.46 0.48
Van-60 60M 0.38 0.46 0.48 0.50 0.51 0.52
Van-60 60M 0.36 0.42 0.44 0.45 0.45 0.46
Table S6: Top-$K$ accuracy on subseta of 500 routes from test set-n1.
SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.402 0.476 0.522 0.550 0.564 0.582
6x4-12x4b 18M 0.446 0.552 0.598 0.616 0.624 0.640
12x3-12x3 19M 0.452 0.570 0.596 0.618 0.626 0.636
12x4-12x4 22M 0.452 0.552 0.588 0.610 0.620 0.642
6x3-6x3 10M 0.292 0.362 0.384 0.388 0.396 0.402
aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
Table S7: Top-$K$ accuracy on subseta of 500 routes from test set-n5.
SM SM
Models Size Steps (Train) (Gen.) Top-1 Top-2 Top-3 Top-4 Top-5 Top-10
Flash 10M 0.374 0.442 0.474 0.484 0.498 0.516
6x4-12x4b 18M 0.400 0.464 0.498 0.520 0.530 0.548
12x3-12x3 19M 0.406 0.472 0.496 0.518 0.524 0.538
12x4-12x4 22M 0.370 0.448 0.472 0.498 0.508 0.526
6x3-6x3 10M 0.222 0.264 0.282 0.288 0.290 0.298
aThe subset was chosen randomly with a seed of 42. Flash results are provided to establish how the performance on these 500 routes relates to the performance on full test set.
bThe model names follow NxF-KxF format, where N is the number of encoder layers, K is the number of decoder layers, and F is the feed-forward multiplier
Table S8: Top-1 Accuracy on n1 routes split by route length
Model 2 3 4 5 6 7 8 9 10
Flash 0.538 0.410 0.323 0.273 0.329 0.308 0.286 0.400 0.000
Explorera 0.480 0.346 0.282 0.266 0.305 0.323 0.286 0.500 0.000
Explorer-XLa,b 0.402 0.304 0.271 0.259 0.296 0.277 0.286 0.600 0.000
Flex (Mono) 0.456 0.318 0.226 0.193 0.183 0.154 0.143 0.300 0.000
Flex (Duo) 0.554 0.426 0.348 0.292 0.362 0.354 0.286 0.500 0.000
Flex (Duo)b 0.471 0.360 0.310 0.278 0.319 0.277 0.286 0.400 0.000
Deepb 0.497 0.395 0.338 0.268 0.272 0.262 0.286 0.600 0.000
Wideb 0.511 0.416 0.344 0.308 0.329 0.292 0.286 0.600 0.000
aSteps not provided during inference.
bSMs not provided during inference.
Table S9: Top-10 Accuracy on n1 routes split by route length
Top10 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.763 0.600 0.465 0.385 0.408 0.400 0.333 0.500 0.000
Explorera 0.692 0.546 0.417 0.373 0.376 0.431 0.333 0.500 0.000
Explorer-XLa,b 0.608 0.463 0.387 0.363 0.390 0.400 0.333 0.600 0.000
Flex (Mono) 0.650 0.474 0.325 0.254 0.286 0.231 0.143 0.300 0.000
Flex (Duo) 0.765 0.610 0.469 0.373 0.408 0.385 0.333 0.600 0.000
Flex (Duo)b 0.594 0.460 0.388 0.336 0.371 0.338 0.333 0.600 0.000
Deepb 0.641 0.519 0.442 0.364 0.362 0.323 0.333 0.600 0.000
Wideb 0.647 0.531 0.434 0.378 0.390 0.354 0.333 0.600 0.000
aSteps not provided during inference.
bSMs not provided during inference.
Table S10: Top-1 Accuracy on n5 routes split by route length
Top1 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.502 0.401 0.337 0.320 0.331 0.328 0.429 0.692 0.333
Explorera 0.458 0.350 0.300 0.289 0.343 0.288 0.286 0.462 0.333
Explorer-XLa,b 0.395 0.317 0.284 0.303 0.313 0.248 0.405 0.692 0.333
Flex (Mono) 0.436 0.324 0.258 0.211 0.191 0.184 0.190 0.154 0.333
Flex (Duo) 0.549 0.416 0.366 0.336 0.352 0.328 0.476 0.615 0.333
Flex (Duo)b 0.465 0.358 0.324 0.312 0.326 0.344 0.429 0.615 0.333
Deepb 0.474 0.384 0.343 0.313 0.338 0.272 0.381 0.385 0.333
Wideb 0.477 0.399 0.352 0.337 0.366 0.304 0.381 0.615 0.333
aSteps not provided during inference.
bSMs not provided during inference.
Table S11: Top-10 Accuracy on n5 routes split by route length
Top10 accuracy on routes with length
Model 2 3 4 5 6 7 8 9 10
Flash 0.712 0.583 0.479 0.431 0.425 0.400 0.429 0.692 0.333
Explorera 0.654 0.529 0.442 0.404 0.434 0.376 0.405 0.692 0.333
Explorer-XLa,b 0.586 0.469 0.416 0.404 0.423 0.384 0.47 0.692 0.333
Flex (Mono) 0.621 0.461 0.357 0.285 0.297 0.272 0.333 0.462 0.333
Flex (Duo) 0.721 0.594 0.489 0.436 0.444 0.416 0.500 0.692 0.333
Flex (Duo)b 0.581 0.462 0.407 0.383 0.421 0.400 0.452 0.692 0.333
Deepb 0.604 0.507 0.451 0.414 0.430 0.368 0.429 0.615 0.333
Wideb 0.616 0.524 0.451 0.420 0.432 0.328 0.476 0.692 0.333
aSteps not provided during inference.
bSMs not provided during inference.